OpenStudy (anonymous):

You have a 2.00 kg block of lead. Lead melts at 327.5°C. CPb = 130.0 J/kg x g C and Hf for lead is 2.04 × 10^4 J/kg. You begin at room temperature (25.0°C). How much heat do you need to transfer to melt all the lead?

OpenStudy (anonymous):

@roadjester

OpenStudy (roadjester):

What do all of your letters mean? I think Hf is latent heat of fusion....

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

yes hf means heat of fusion

OpenStudy (ybarrap):

I'm just learning thermodynamics, but let's see if we can work this out. Is \(C_{Pb}\) the heat capacity for Lead? We probably need to convert C to K for temp. What do you think are the relevant equations? Is it \(U=nC_{Pb}\Delta T\) Joules?

OpenStudy (ybarrap):

*If we only care about differences, we don't need to convert C to K.

OpenStudy (anonymous):

yeah my teacher sees every little detail so she wants every conversion and units to be perfect

OpenStudy (anonymous):

@roadjester

OpenStudy (roadjester):

I haven't done thermodynamics in a while since it was two semesters ago i think

OpenStudy (roadjester):

You're going to need two equations for this. Q=mc\(\Delta\)T and Q=mL Where the first one is when you are changing temperature and the second is changing between states.

OpenStudy (roadjester):

Ok, what the heck is CPb?

OpenStudy (roadjester):

I'm guessing Pb is lead but what is CPb?

OpenStudy (anonymous):

i think the Cpb means the heat capacity for lead like @ybarrap said

OpenStudy (roadjester):

m=2kg \(\Large c_{Pb}=130...\) \(\color {red} {what \space\space the \space\space heck\space\space is\space\space up\space\space with\space\space the\space\space units?}\) \(\large L_f(Latent \space heat\space of\space \space fusion)\)=2.04 X10^4 J/kg T_i =25 degrees C

OpenStudy (roadjester):

you've got Joules per, gram, kilogram, AND C??? four units?

OpenStudy (anonymous):

yep i think so I've got : 2.00 kg block of lead lead melts at 327.5 C Cpb = 130.0 J/kg x g degrees C Hf = 2.04 x 10^4 J/kg for lead

OpenStudy (roadjester):

there is no way you can have two units of mass

OpenStudy (anonymous):

start at room temperature of 25 degrees celsius then find how much heat is needed to transfer to melt the lead

OpenStudy (roadjester):

I know that, I've taken thermophysics

OpenStudy (roadjester):

but your units are screwy

OpenStudy (anonymous):

yeah it looks weird maybe they messed up on it

OpenStudy (anonymous):

i think the g needs to be taken out and just remain with kg

OpenStudy (roadjester):

either g or kg, can't have both

OpenStudy (ybarrap):

$$ \large{ U_2-U_1=m\cdot c \cdot (T_2-T_1)\\ =2~kg\times 130.0~ J/kg\cdot K \times(327.5-25) \\ =78,650~J\\ (Q_{lead})_{in}=h_f \times m_{Pb}=2.04 × 10^4~J/kg\times2kg=40,800 J\\ } $$ http://hyperphysics.phy-astr.gsu.edu/hbase/tables/sphtt.html The change in internal energy in changing the temperature of the lead from 25 to 327.5 C is 78,650 J. However, only 40,800 J is required to melt all the lead. The remaining energy will be stored in the liquid lead.

OpenStudy (ybarrap):

I may be wrong here. It might be that 78,650 J is required to melt the lead because you need to get to the temp of 327.5 C to melt. Once melting is complete, 40,800 J would have been released through latent heat of fusion and the remaining energy stored in the liquid.

OpenStudy (anonymous):

yes they ask me for the amount of heat needed to be transferred to melt the entire lead

OpenStudy (anonymous):

but i think what you got was correct

OpenStudy (ybarrap):

I think my 1st answer is correct, 40,800 J. I just checked here and get approximately the same answer -- I also validated for gold - http://wiki.answers.com/Q/How_much_energy_is_it_required_to_melt_2_kg_of_gold#slide=2

OpenStudy (anonymous):

for where u put kg is it a specfici amount or just kg?

OpenStudy (ybarrap):

I used this latent heat of fusion chart for other elements - http://www.engineeringtoolbox.com/fusion-heat-metals-d_1266.html BTW, your heat of fusion for lead is 2.04 × 10^4 J/kg while the site above shows it to be 2.3 × 10^4 J/kg, which really doesn't matter that much. But when I used the 2.3 × 10^4 J/kg, I got the exact same answer as wiki.answers.com for lead, 36.8 kJ, versus my answer of 40.8 kJ.

OpenStudy (ybarrap):

I don't understand your question

OpenStudy (anonymous):

does kg stand for a specific number or is it just supposed to be written there?

OpenStudy (ybarrap):

I just wrote the units (for example, kg) so you are aware what the number preceding it is supposed to represent.

OpenStudy (ybarrap):

just helps with dimension analysis, to validate that units cancel appropriately

OpenStudy (ybarrap):

The only problem I have with this solution, if it is in fact 40.8 kJ is that this did not require us to use temperature at all. Perhaps if we found that there was not enough energy generated by the change in internal temperature, then we could not say that all the lead was melted. So in a sense, we DID use that information.

OpenStudy (anonymous):

so was 78,650 supposed to be part of the answer?

OpenStudy (ybarrap):

No. I think that if THAT number, taking into account temperature change (i.e. the change in internal energy) were less than 40.8 kJ, then we could not say that ALL of the lead would have melted.

OpenStudy (ybarrap):

It IS part of the answer to prove that the change in internal energy was higher than 40.8 kJ, which was required to melt ALL the lead.

OpenStudy (anonymous):

so 78,650 is the heat required to be transferred to melt all the lead?

OpenStudy (ybarrap):

No, the heat required to melt ALL the lead is just 40.8kJ. The other number was the ACTUAL heat that was transferred. So the lead received more heat than was required to melt it.

OpenStudy (ybarrap):

Remember, you can heat a solid until it melts. And then you can add more heat after it is melted. This additional heat will just be stored in the liquid until there is enough heat to turn the liquid into a gas. That's what was done here -- but we did not check if it started turning into a gas.

OpenStudy (anonymous):

oh ok yeah it just asks for the amount of heat needed to be transferred

OpenStudy (anonymous):

so my answer would be 78,650?

OpenStudy (ybarrap):

No, the heat required to melt ALL the lead is just 40.8kJ.

OpenStudy (anonymous):

so 40,800 kJ is final answer to "How much heat must you transfer to melt all the lead?"

OpenStudy (anonymous):

40,800 J because kg's cancel out* sorry

OpenStudy (ybarrap):

Yes.